### v2

# TextOCVP: Object-Centric Video Prediction with Language Guidance

Abstract
Understanding and forecasting future scene states is critical for autonomous agents to plan and act effectively in complex environments. Object-centric models, with structured latent spaces, have shown promise in modeling object dynamics and predicting future scene states, but often struggle to scale beyond simple synthetic datasets and to integrate external guidance, limiting their applicability in robotics. To address these limitations, we propose TextOCVP, an object-centric model for video prediction guided by textual descriptions. TextOCVP parses an observed scene into object representations, called slots, and utilizes a text-conditioned transformer predictor to forecast future object states and video frames. Our approach jointly models object dynamics and interactions while incorporating textual guidance, enabling accurate and controllable predictions. TextOCVP’s structured latent space offers a more precise control of the forecasting process, outperforming several video prediction baselines on two datasets. Additionally, we show that structured object-centric representations provide superior robustness to novel scene configurations, as well as improved controllability and interpretability, enabling more precise and understandable predictions. Videos and code are available at https://play-slot.github.io/TextOCVP.
1 Introduction
Understanding and reasoning about the environment is essential for enabling autonomous systems to better comprehend their surroundings, predict future events, and adapt their actions accordingly. Humans achieve these capabilities by perceiving the environment as a structured composition of individual objects that interact and evolve dynamically over time (Kahneman, Treisman, and Gibbs 1992). Neural networks equipped with such compositional inductive biases have shown the ability to learn structured object-centric representations of the world, which enable desirable properties, such as out-of-distribution generalization (Dittadi et al. 2022), compositionality (Greff, Van Steenkiste, and Schmidhuber 2020), or sample efficiency (Mosbach et al. 2025).
Recent advances in unsupervised object-centric representation learning have progressed from extracting object representations in synthetic images (Locatello et al. 2020; Lin et al. 2020) to modeling objects in video (Kipf et al. 2022; Singh, Wu, and Ahn 2022) and scaling to real-world scenes (Seitzer et al. 2023; Zadaianchuk, Seitzer, and Martius 2024). These developments have enabled object-level dynamics modeling for future prediction and planning. Notably, approaches like SlotFormer (Wu et al. 2023) or OCVP (Villar-Corrales, Wahdan, and Behnke 2023) introduced object-centric prediction models that explicitly model spatio-temporal relationships between objects, shifting away from image-level approaches that ignore scene compositionality. Despite these advancements, current object-centric methods struggle with complex object appearances and dynamics, and lack mechanisms to incorporate external guidance, thus limiting their scalability and broader applicability.
Refer to captiona) Object-Centric ParsingLearned ObjectsRefer to captionb) Text-Conditioned Video Prediction
‘Put the green block
in the cyan bowl.’
‘Put the green block
in the red bowl.’
Figure 1: Overview of TextOCVP. (a) Our model parses a reference frame $\textbf{X}_{1}$ into its object components $\textbf{S}_{1}$. (b) Our TextOCVP predictor jointly models object dynamics and interactions guided by text, generating future object states and frames that align with the provided textual instructions.
To address these challenges, we propose TextOCVP, a novel object-centric model for video prediction guided by textual instructions, illustrated in Fig. 1. Given a reference image and text instruction, TextOCVP extracts object representations and predicts their evolution using a text-conditioned object-centric transformer. This predictor forecasts future object states by explicitly modeling their dynamics and interactions over time, while integrating textual information via a text-to-slot attention mechanism. By jointly modeling spatio-temporal object relationships and incorporating textual guidance, TextOCVP predicts future object states and frames aligned with the input instruction.
We evaluate our approach for text-conditioned video prediction through extensive experiments on two distinct datasets. Our results show that TextOCVP outperforms existing text-conditioned video prediction methods by effectively leveraging structured object-centric representations, particularly on scenes featuring multiple moving objects.
To verify the effectiveness of object-centric representations for text-guided video prediction, we conduct an in-depth analysis of TextOCVP. Specifically, we demonstrate that its structured latent space enables accurate and controllable video prediction by aligning language instructions with the corresponding objects, outperforming baselines that rely on holistic scene representations. Beyond generation quality, TextOCVP offers improved interpretability and exhibits strong robustness to novel scene configurations, including varying numbers of objects or previously unseen colors.
In summary, our contributions are as follows:
  • We propose TextOCVP, a text-guided video prediction model, featuring a text-conditioned object-centric predictor that integrates textual guidance into the prediction process via a text-to-slot attention mechanism.
  • Through extensive evaluations, we show that TextOCVP outperforms existing text-conditioned video prediction models by leveraging object-centric representations.
  • We demonstrate that TextOCVP is controllable, seamlessly adapting to diverse textual instructions, while exhibiting interpretability and robustness to novel setups.
2 Related Work
2.1 Object-Centric Learning
Representation learning, the ability to extract meaningful features from data, often improves model performance by enhancing its understanding of the input space (Bengio, Courville, and Vincent 2013). Object-centric representation methods aim to parse an image or video into a set of object components in an unsupervised manner. These objects are typically represented as unconstrained embeddings (called slots) (Locatello et al. 2020; Kipf et al. 2022; Singh, Wu, and Ahn 2022), patch-based representations (Lin et al. 2020), factored latent vectors (Greff et al. 2019), or explicit object prototypes (Villar-Corrales and Behnke 2022; Monnier, Groueix, and Aubry 2020). These methods have demonstrated promise in learning object representations across diverse domains, ranging from synthetic images (Locatello et al. 2020; Lin et al. 2020) to videos (Kipf et al. 2022; Singh, Wu, and Ahn 2022; Elsayed et al. 2022), and real-world scenes (Seitzer et al. 2023; Zadaianchuk, Seitzer, and Martius 2024). The learned object representations benefit downstream tasks, such as reinforcement learning (Mosbach et al. 2025) or visual-question answering (Wu et al. 2023).
2.2 Video Prediction
Video prediction (VP) is the task of forecasting the upcoming $T$ video frames conditioned on the preceding $C$ seed frames (Oprea et al. 2020). Several methods have been proposed to address this challenge, leveraging 2D convolutions (Gao et al. 2022; Chiu, Adeli, and Niebles 2020), 3D convolutions (Tulyakov et al. 2018), recurrent neural networks (RNNs) (Denton and Fergus 2018; Villar-Corrales et al. 2022; Wang et al. 2022), Transformers (Rakhimov et al. 2021; Ye and Bilodeau 2022), or diffusion models (Höppe et al. 2022; Ho et al. 2022).
Object-Centric Video Prediction
Object-centric VP presents a structured approach that explicitly models the dynamics and interactions of individual objects to forecast future video frames. These methods typically involve three main steps: parsing seed frames into object representations, forecasting future object states using a dynamics model, and rendering video frames from the predicted object representations. Various approaches have addressed this task using different architectural priors, such as RNNs (Creswell et al. 2021; Nguyen et al. 2024) or transformers (Wu, Yoon, and Ahn 2021; Wu et al. 2023; Villar-Corrales, Wahdan, and Behnke 2023). Despite promising results, these models are limited to simple deterministic datasets or rely on action-conditioning (Mosbach et al. 2025; Villar-Corrales and Behnke 2025). In contrast, our model forecasts future frames conditioned on past object slots and text descriptions.
Text-Conditioned Video Prediction
Text-conditioned VP models leverage text descriptions to provide appearance, motion and action cues that guide the generation of future frames. This task was first proposed by Hu, Luo, and Chen (2022), who utilized a VQ-VAE to encode images into visual token representations, and modeled the scene dynamics with an axial transformer to jointly process visual tokens with text descriptions. Similarly, approaches like TVP (Song et al. 2024) and MMVG (Fu et al. 2023) address this task using RNNs or masked transformers, respectively. More recently, several methods leverage diffusion models for text-guided VP (Blattmann et al. 2023; Ni et al. 2023; Chen et al. 2024; Xing et al. 2024; Chen et al. 2023; Gu et al. 2024). These approaches encode video frames into discrete token sequences via pretrained quantized autoencoders, and leverage pretrained language models to guide the diffusion-based generation with text features. To further improve temporal coherence and semantic alignment, several diffusion-based methods leverage specialized conditioning strategies (Xing et al. 2024; Chen et al. 2023) or attention mechanisms (Gu et al. 2024), among others. While effective, these models operate on holistic or spatial representations, and require large-scale compute and data. In contrast, TextOCVP adopts a more structured and efficient approach, explicitly modeling object dynamics using slot representations, where each object in the scene is represented by a distinct embedding.
Concurrently with our work, Wang et al. (2025) and Jeong et al. (2025) combine object-centric learning with autoregressive diffusion models and transformers for text-guided video prediction on simple synthetic datasets. In contrast, we model object dynamics with an autoregressive transformer, evaluate on more complex robotics simulations, and perform an in-depth analysis of the properties of object-centric representations for text-conditioned video prediction.
3 Method
We propose TextOCVP, a novel object-centric model for text-conditioned video prediction. Given an initial reference image $\textbf{X}_{1}$ and a text caption $\mathcal{C}$, TextOCVP generates the $T$ subsequent video frames $\hat{\mathbf{X}}_{2:{T+1}}$, which maintain a similar appearance and structural composition as the reference image, and follow the motion described in the text caption.
TextOCVP, which is illustrated in Fig. 2, implements an object-centric approach, in which the reference frame $\textbf{X}_{1}$ is first decomposed with a scene parsing module (Sec. 3.1) into a set of $N_{\mathbf{S}}$ $D$-dimensional object representations called slots $\textbf{S}_{1}\in{\mathbb{R}}^{N_{\mathbf{S}}\times D}$, where each slot represents a single object in the image. The object slots are fed to a text-conditioned transformer predictor (Sec. 3.2), which jointly models their spatio-temporal relations, and incorporates the textual information from the caption $\mathcal{C}$ as guidance for predicting the future object slots $\hat{\mathbf{S}}_{2:T+1}$. Finally, the predicted slots are decoded to render future video frames (Sec. 3.3).
We propose two different TextOCVP variants, which differ in the underlying object-centric decomposition modules. Specifically, $\text{TextOCVP}_{\text{SAVi}}$ leverages SAVi (Kipf et al. 2022), whereas $\text{TextOCVP}_{\text{DINO}}$ extends the recent DINOSAUR (Seitzer et al. 2023) framework for recursive object-centric video decomposition and video rendering.
3.1 Scene Parsing
The scene parsing module decomposes a video sequence $\mathbf{X}_{1:{\tau}}$ into a set of permutation-invariant object representations called slots $\mathbf{S}_{1:\tau}=(\textbf{S}_{1},...,\textbf{S}_{\tau})$, with $\textbf{S}_{t}=(\textbf{s}_{t}^{1},...,\textbf{s}_{t}^{N_{\mathbf{S}}})$, where each slot $\textbf{s}\in{\mathbb{R}}^{D}$ represents a single object. For scene parsing, we adopt the recursive object-centric video decomposition framework from Kipf et al. (2022).
At time step $t$, the corresponding input frame $\textbf{X}_{t}$ is encoded with a feature extractor module into a set of $D_{h}$-dimensional feature maps $\textbf{h}_{t}\in{\mathbb{R}}^{L\times D_{h}}$ representing $L$ spatial locations. The feature extractor is a convolutional neural network in our $\text{TextOCVP}_{\text{SAVi}}$ variant and a DINO-pretrained vision transformer (Caron et al. 2021) in $\text{TextOCVP}_{\text{DINO}}$. These feature maps are processed with Slot Attention (Locatello et al. 2020), which updates the previous slots $\textbf{S}_{t-1}$ based on visual features from the current frame following an iterative attention mechanism. Namely, Slot Attention performs cross-attention, with the attention weights normalized over the slot dimension, thus encouraging the slots to compete to represent parts of the input. It then updates the slots using a Gated Recurrent Unit (Cho et al. 2014) (GRU). Formally, Slot Attention updates the previous slots $\textbf{S}_{t-1}$ by:
$\textbf{A}=\operatorname*{softmax}_{N_{\mathbf{S}}}\left(\frac{q(\textbf{S}_{t-1})k(\textbf{h}_{t})^{T}}{\sqrt{D}}\right)\in{\mathbb{R}}^{N_{\mathbf{S}}\times L},$ (1)
$\textbf{S}_{t}=\text{GRU}(W_{t}v(\textbf{h}_{t}),\textbf{S}_{t-1})\;\text{with}\;W_{i,j}=\frac{\textbf{A}_{i,j}}{\sum_{l=1}^{L}\textbf{A}_{i,l}},$ (2)
where $k,q$ and $v$ are learned linear projections that map input features and slots into a common dimension. The output of this module is a set of slots $\textbf{S}_{t}$ that represents the objects of the input frame.
Figure 2: Overview of TextOCVP. TextOCVP parses the reference frame $\textbf{X}_{1}$ into object representations $\textbf{S}_{1}$. The text-conditioned object-centric predictor models object dynamics and interactions, incorporating information from the description $\mathbf{C}$ to predict future object states $\hat{\mathbf{S}}_{2:T+1}$, which can be used to render future video frames $\hat{\mathbf{X}}_{2:{T+1}}$.
3.2 Text-Conditioned Object-Centric Predictor
Our proposed text-conditioned predictor module, depicted in Fig. 3, autoregressively forecasts future object states conditioned on the object slots from the reference frame $\textbf{S}_{1}$ and a textual description $\mathcal{C}$.
To condition the prediction process, the text description $\mathcal{C}$ is encoded into text token embeddings $\mathbf{C}$ employing an encoder-only transformer module. We experiment with different variants of this module, including a vanilla transformer encoder (Vaswani et al. 2017) or a pretrained T5 (Raffel et al. 2020) module.
At time step $t$, the predictor receives as input the corresponding text embeddings $\mathbf{C}$, as well as the previous object slots $\mathbf{S}_{1:t}$, which are initially mapped via an MLP to the predictor token dimensionality. Additionally, these tokens are augmented with a temporal positional encoding, which applies the same sinusoidal positional embedding to all tokens from the same time step, thus preserving the inherent permutation-equivariance of the objects.
Each layer of our predictor module mirrors the transformer decoder architecture (Vaswani et al. 2017). First, a self-attention layer enables every slot to attend to all other object representations in the sequence, modeling the spatio-temporal relations between objects. Subsequently, a text-to-slot cross-attention layer enhances the slot representations by incorporating relevant features from the text embeddings, guiding the prediction process to align with the motion and dynamics described in the textual caption. Finally, an MLP is applied for each token. This process is repeated in every predictor layer, resulting in the predicted object slots of the subsequent time step $\hat{\textbf{S}}_{t+1}$. Furthermore, we apply a residual connection from $\textbf{S}_{t}$ to $\hat{\textbf{S}}_{t+1}$, which improves the prediction temporal consistency. This process is repeated autoregressively to obtain slot predictions for $T$ subsequent time steps.
3.3 Video Rendering
The video rendering module decodes the predicted slots $\hat{\textbf{S}}_{t}$ to render the corresponding video frame $\hat{\textbf{X}}_{t}$. We leverage two variants of the video rendering module, for our $\text{TextOCVP}_{\text{SAVi}}$ and $\text{TextOCVP}_{\text{DINO}}$ variants, respectively.
$\text{TextOCVP}_{\text{SAVi}}$ Decoder
This variant independently decodes each slot in $\hat{\textbf{S}}_{t}$ with a CNN-based Spatial Broadcast Decoder (Watters et al. 2019), rendering an object image $\textbf{o}_{t}^{n}$ and mask $\textbf{m}_{t}^{n}$ for each slot $\textbf{s}_{t}^{n}$. The object masks are normalized across the slot dimension, and the representations are combined via a weighted sum to render video frames:
$\hat{\textbf{X}}_{t}=\sum_{n=1}^{N_{\mathbf{S}}}\textbf{o}_{t}^{n}\cdot\tilde{\textbf{m}}_{t}^{n}\;\;\;\text{with}\;\;\;\tilde{\textbf{m}}_{t}^{n}=\operatorname*{softmax}_{N_{\mathbf{S}}}(\textbf{m}_{t}^{n}).$ (3)
Figure 3: Overview of our text-conditioned object-centric video predictor. Our predictor employs a transformer-based architecture to model the spatio-temporal dynamics of object slots via self-attention, while incorporating textual guidance through a text-to-slot cross-attention mechanism.
$\text{TextOCVP}_{\text{DINO}}$ Decoder
This decoder variant decodes the object slots in two distinct stages. First, following DINOSAUR (Seitzer et al. 2023), an MLP-based Spatial Broadcast Decoder (Watters et al. 2019) is used to generate object features along with their corresponding masks. Similar to the $\text{TextOCVP}_{\text{SAVi}}$ decoder, the object masks are normalized and combined with the object features in order to reconstruct the encoded features $\hat{\textbf{h}}_{t}\in{\mathbb{R}}^{L\times D_{h}}$. In the second stage, the reconstructed features $\hat{\textbf{h}}_{t}$ are arranged into a grid format and processed with a CNN decoder to generate the corresponding video frame $\hat{\textbf{X}}_{t}$.
3.4 Training and Inference
Our proposed TextOCVP is trained in two different stages.
Object-Centric Learning
We first train the scene parsing and video rendering modules for parsing video frames into object-centric representations by minimizing a reconstruction loss. In the $\text{TextOCVP}_{\text{SAVi}}$ variant ($\mathcal{L}_{\text{SAVi}}$), these modules are trained by reconstructing the input images, whereas in $\text{TextOCVP}_{\text{DINO}}$ ($\mathcal{L}_{\text{DINO}}$) they are trained by jointly minimizing an image and a feature reconstruction loss:
[…]
Predictor Training
Given the pretrained scene parsing and rendering modules, we train our TextOCVP predictor for text-conditioned video prediction using a dataset containing paired videos and text descriptions. Namely, given the object representations from a reference frame $\textbf{S}_{1}$ and the textual embeddings $\mathbf{C}$, our TextOCVP predictor forecasts the subsequent object slots $\hat{\textbf{S}}_{2}$, which are decoded into a predicted video frame $\hat{\textbf{X}}_{2}$. This process is repeated autoregressively, i.e. the predicted slots are appended to the input in the next time step, in order to generate the set of slots for the subsequent $T$ time steps. This autoregressive training, in contrast to teacher forcing, enforces our predictor to operate with imperfect inputs, leading to better modeling of long-term dynamics at inference time.
Our predictor is trained by minimizing the following loss:
$\mathcal{L}_{\text{TextOCVP}}=\frac{1}{T}\sum_{t=2}^{T+1}\left(\lambda_{\text{Img}}\mathcal{L}_{\text{Img}}+\lambda_{\text{Slot}}\mathcal{L}_{\text{Slot}}\right),$ (6)
[…]
where $\mathcal{L}_{\text{Img}}$ measures the future frame prediction error, and $\mathcal{L}_{\text{Slot}}$ enforces the alignment of the predicted object slots with the actual inferred object-centric representations.
Inference
At inference time, TextOCVP receives as input a single reference frame and a language instruction. Our model parses the seed frame into object slots and autoregressively predicts future object states and video frames conditioned on the given textual description. By modifying the language instruction, TextOCVP can generate a new sequence continuation that performs the specified task while preserving a consistent scene composition.
4 Experiments
4.1 Experimental Setup
Datasets
We evaluate TextOCVP for text-conditioned video prediction on the CATER and CLIPort datasets:
CATER (Girdhar and Ramanan 2020) is a dataset of long video sequences showing 3D objects in motion, each accompanied by a descriptive caption. We used the CATER-hard variant (Hu, Luo, and Chen 2022), containing 30,000 sequences with $64\times 64$ frames featuring between two and eight objects, two of which follow the scripted motion.
CLIPort (Shridhar, Manuelli, and Fox 2022) is a robot manipulation dataset featuring video-caption pairs. We employ 21,000 $336\times 336$ sequences of the Put-Block-In-Bowl task, where the caption describes placing a block into a bowl.
Baselines
[…]
Refer to captiont=1t=122101020203030MAGE TextOCVP
‘the medium green metal sphere is sliding
to (2, 1). the small brown metal cube
is picked up and placed to (-3, 1)’
Figure 4: Text-guided video prediction result on CATER. Top row shows ground truth. TextOCVP predicts sharp and accurate frames, whereas MAGE blurs and misses objects.
Implementation Details
All models are implemented in PyTorch and trained on a single NVIDIA A6000 (48Gb) GPU. $\text{TextOCVP}_{\text{SAVi}}$ closely follows Kipf et al. (2022) for scene parsing and video rendering. $\text{TextOCVP}_{\text{DINO}}$ uses DINOv2 (Oquab et al. 2024) as image encoder, a four-layer MLP Spatial-Broadcast Decoder (Watters et al. 2019) to decode the slots into object features and masks, and a CNN decoder to map the reconstructed scene features back to images. On CATER, we use the $\text{TextOCVP}_{\text{SAVi}}$ variant with eight 128-dimensional object slots, whereas on CLIPort we employ the $\text{TextOCVP}_{\text{DINO}}$ variant with ten 128-dimensional slots. Our predictor module is an eight-layer transformer with 512-dimensional tokens, eight attention heads, and a hidden dimension of 1024.
Further experiment details are provided in Appendix B-E.
4.2 Results
CATER Results
On CATER, we train the models to predict nine future frames given a single reference frame and a text caption. In Table 1(a) we report quantitative evaluations on CATER using the same setting as in training, i.e. predicting $T=9$ frames, as well as when predicting $T=19$ future frames. In both settings, TextOCVP outperforms all other models, demonstrating superior perceptual quality.
Fig. 4 qualitatively compares TextOCVP and MAGE. Our proposed method generates a sequence that is closely aligned to the ground-truth, whereas MAGE predictions feature multiple errors and artifacts, including missing objects, blurry contours, and significant changes on object shapes. This highlights the effectiveness of object-centric representations to accurately represent and model object dynamics.
[…]
(a) CATER Evaluation
CATER1→9 CATER1→19
[…]
OCVP 0.874 0.078 4.16 0.854 0.101 8.08
Non-OC 0.874 0.092 3.04 0.849 0.112 8.62
SEER 0.723 0.245 11.23 0.535 0.299 17.29
MAGE 0.877 0.108 3.46 0.871 0.111 5.88
TextOCVP 0.922 0.036 2.16 0.902 0.044 5.09
(b) CLIPort Evaluation
CLIPort1→9 CLIPort1→19
[…]
Non-OC 0.901 0.184 8.13 0.872 0.210 13.23
SEER 0.887 0.141 6.80 0.622 0.331 8.29
$\text{MAGE}_{\text{DINO}}$ 0.940 0.064 2.11 0.931 0.075 2.59
TextOCVP 0.950 0.062 1.36 0.931 0.078 2.23
Table 2: Ablation studies. We ablate different architectural design choices in our proposed TextOCVP model.
(a) Effect of number of layers in predictor.
CATER1→9
[…]
2 0.908 0.045 2.26
4 0.911 0.043 2.20
8 0.922 0.036 2.16
(b) Evaluation of different text encoders on CATER and CLIPort.
CATER1→19 CLIPort1→19
[…]
Custom TF 0.903 0.045 5.39 - - -
Frozen T5 0.902 0.044 5.09 0.931 0.078 2.23
FT T5 0.901 0.043 5.14 0.928 0.082 2.67
(c) Effect of residual connection in predictor.
CLIPort1→9
[…]
No 0.946 0.066 1.58
Yes 0.950 0.062 1.36
(d) Effect of the number of slots $N_{\mathbf{S}}$.
CLIPort1→9
[…]
8 0.946 0.079 5.62
10 0.950 0.062 1.36
CLIPort Results
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
‘put the blue block in the gray bowl
[…]
TextOCVPt=1t=12020303040405050Missing
[…]
Nevertheless, we identify that TextOCVP often generates visual artifacts and lacks fine-grained textures, which can affect its quantitative scores despite accurate predictions.
Further figures and results are provided in Appendix F.
4.3 Model Analysis
Ablation Studies
We perform several ablation studies to support and validate the architectural choices of our model components and their impact on TextOCVP’s video prediction performance. The results are presented in Table 2(d).
Number of Layers (Table 2(a))   TextOCVP performs best with $N_{\text{P}}=8$ predictor layers, outperforming shallower variants. We did not explore scaling beyond eight layers due to the significant increase in the number of parameters.
Residual Connection (Table 2(c))   Incorporating a residual connection in the predictor, $\hat{\textbf{S}}_{t+1}:=\hat{\textbf{S}}_{t+1}+\textbf{S}_{t}$, enhances the model’s performance by improving the temporal consistency of the predicted slots.
Text Encoder (Table 2(b))   We evaluate the performance of different text encoders, including a vanilla transformer, a frozen T5 encoder, and a fine-tuned (FT) T5. A frozen T5 module led to the best TextOCVP performance on the CLIPort dataset. Although all encoders led to similar performance on the CATER dataset, we opted for the frozen T5 to avoid unnecessarily increasing the model’s parameter count.
Number of Slots (Table 2(d))   Using $N_{\mathbf{S}}=10$ slots resulted in better performance for TextOCVP on CLIPort, even though each scene can be described with eight slots (six objects, robot arm and background). This observation suggests that additional slots can be used as registers in order to help with internal computations (Darcet et al. 2024).
Table 3: Quantitative evaluation on a CLIPort test set with unseen colors. We report the video prediction performance and its drop relative to the known-color dataset.
CLIPort1→9 CLIPort1→19
[…]
$\text{MAGE}_{\text{DINO}}$ 0.935 -0.5% 0.076 -19% 0.924 -0.7% 0.087 -16%
TextOCVP 0.946-0.4% 0.066-6.4% 0.927-0.4% 0.083-6.4%
Model Robustness
‘the large purple rubber cone is picked up
and placed to (2, 3). the small gold metal
snitch is picked up and placed to (-1, 1).’
Original
Caption
a)
‘the medium cyan rubber sphere is picked up
and placed to (-2, 2). the medium purple
metal cone is sliding to (2, 3).’
Changed
Objects
& Actions
b)
‘the large purple rubber cone is sliding to
(-1, 1). the small gold metal snitch is
rotating. the medium purple metal cone
is picked up and placed to (-1, -3).’
More
Moving
Objects
c)
t=1t=12215153030
Figure 6: Evaluation of TextOCVP’s controllability on CATER. Top row shows the ground truth sequence. TextOCVP enables predicting different sequence continuations conditioned on distinct text instructions.
We evaluate the generalization ability of TextOCVP and $\text{MAGE}_{\text{DINO}}$ on a CLIPort evaluation set featuring color variations in the text instructions that were not encountered during training. This setup assesses the models’ ability to generalize to novel scene configurations and adapt to unseen text-visual combinations.
Table 3 presents the results on this evaluation set with unseen colors. We report both the video prediction performance as well as the performance drop relative to the evaluation set with known colors. Our model consistently outperforms $\text{MAGE}_{\text{DINO}}$ for both prediction horizons. Most notably, TextOCVP demonstrates significantly higher robustness to objects with previously unseen colors compared to the baseline, showing a much smaller drop in performance for the perceptual LPIPS metric. These results highlight the advantages of structured object-centric representations, which provide greater generalization and robustness in video prediction—particularly in contrast to holistic scene representations that struggle with novel scene compositions.
Controllability
A key objective of text-driven video prediction is to provide fine-grained control over the prediction process, which is enabled through language instructions describing the objects in the scene and their expected motion.
In Fig. 6, we qualitatively showcase the level of control that TextOCVP provides over its generations on the CATER dataset. We qualitatively evaluate this ability by generating multiple video continuations from the same reference frame while varying the language instruction. These variations include changes to both the target objects and their actions, as well as providing instructions that include a number of actions beyond what the model was trained on. As shown in Fig. 6, TextOCVP successfully identifies the relevant objects and executes the actions described in the text instructions. Notably, our model distinguishes between two nearly identical purple cones, despite the identical shapes and color, and correctly generates a sequence with the specified motion.
These results show that key object attributes, such as size or color, are effectively captured through the text-to-slot attention mechanism, enabling the predictor to accurately forecast the motion of each individual object. This highlights the benefits of combining object-centric representations with text guidance for fine-grained, controllable video prediction.
Interpretability
Refer to captiont=1t=1Refer to caption55Refer to caption1010Refer to caption2020Refer to caption3030
‘the cone is sliding to (-3, -3).
the snitch is picked up and placed to (-1, 2 ).’
Refer to captionRefer to caption1.00.50.0Refer to caption
Slot 3
the
cone
is
sliding
to
-3
-3
the
snitch
is
picked
up
and
placed
to
-1
2
Refer to captionRefer to caption1.00.50.0Refer to caption
Slot 4
the
cone
is
sliding
to
-3
-3
the
snitch
is
picked
up
and
placed
to
-1
2
Refer to captionRefer to caption1.00.50.0Refer to caption
Slot 6
the
cone
is
sliding
to
-3
-3
the
snitch
is
picked
up
and
placed
to
-1
2
Figure 7: Text-to-slot attention weights for different objects in a video. Slots representing objects described in the caption focus on relevant text tokens, such as target locations.
In Fig. 7, we visualize the text-to-slot attention weights—averaged across attention heads—for different slots in a CATER sequence, revealing how textual information guides model predictions. We observe that slots corresponding to objects described in the textual instruction focus on relevant text tokens, such as target coordinate locations. These results demonstrate that the text-to-slot attention mechanism effectively aligns the textual information with the object-centric representations, enabling the model to generate accurate video sequences based on the given text instructions, while providing superior interpretability.
5 Conclusion
We presented TextOCVP, a novel object-centric model for text-conditioned video prediction. Given a single input image and a text description, TextOCVP generates frames by parsing the scene into object representations and modeling their dynamics conditioned on the textual instruction. This is accomplished through a text-conditioned object-centric transformer, which predicts future object states by modeling the spatio-temporal relationships between objects while incorporating textual guidance. Through extensive evaluations, we demonstrated that TextOCVP outperforms existing models for text-driven video prediction from a single frame, highlighting our model’s ability to predict over long prediction horizons and adapt its predictions based on the provided textual description. Moreover, we validated our architectural choices through ablation studies, highlighting the importance of combining textual and object-centric information, as well as demonstrating TextOCVP’s robustness and interpretability. With its structured latent space and superior controllability, TextOCVP offers potential for applications in robotics, where it can serve as a controllable world model, enabling efficient planning, reasoning, and decision-making.
Acknowledgment
This work was funded by grant BE 2556/16-2 (Research Unit FOR 2535 Anticipating Human Behavior) of the German Research Foundation (DFG)
Appendix A Limitations and Future Work
A.1 Limitations
While TextOCVP demonstrates promising results for text-guided object-centric video prediction, it presents some limitations that we plan to address in future work:
Prediction Artifacts
TextOCVP occasionally generates artifacts in the predicted frames, such as blurriness, inconsistent object appearances, lack of textured details, or visual artifacts in the background. We believe that these limitations stem from the video rendering module, which might lack the representational power to reconstruct precise image details from the object-centric latent space representation.
Limited Temporal Consistency
While TextOCVP produces plausible predictions, we observe that its temporal consistency can degrade when forecasting for long prediction horizons ($T>30$), occasionally resulting in object jittering or instability. We attribute this limitation to the fact that TextOCVP is trained to predict only up to nine future frames ($T=9$) and only optimizing reconstruction losses, which do not penalize such temporal inconsistencies.
A.2 Future Work
In future work we aim to address these limitations. Our TextOCVP model is designed with a modular architecture, enabling for both improvements and flexible swapping of components—such as the parsing, predictor, or video rendering modules—to seamlessly improve the entire pipeline.
To address the prediction artifacts, we plan to extend our TextOCVP framework with more powerful decoder modules, such as autoregressive transformers (Singh, Wu, and Ahn 2022) or diffusion models (Jiang et al. 2023), as well as scale our predictor module. Furthermore, we plan to incorporate temporal discriminators (Clark, Donahue, and Simonyan 2020) to improve the temporal consistency of the predicted video frames.
We believe that exploring these architectural modifications can overcome the aforementioned limitations and will enable the scaling of TextOCVP to real-world robotic environments.
Appendix B Implementation Details
We employ $\text{TextOCVP}_{\text{SAVi}}$ for the experiments on CATER and $\text{TextOCVP}_{\text{DINO}}$ for experiments on CLIPort. Below we discuss the implementation details for each of these variants.
B.1 $\text{TextOCVP}_{\text{DINO}}$
The $\text{TextOCVP}_{\text{DINO}}$ variant consists of our proposed text-conditioned predictor module and an object-centric decomposition module that extends the DINOSAUR (Seitzer et al. 2023) framework for recursive object-centric video decomposition and video rendering.
Text-Conditioned Predictor
The predictor is composed of $N_{\text{P}}=8$ identical layers, each containing 8-head attention mechanisms and an MLP with a single hidden layer of dimension 1024 and a ReLU activation function. Furthermore, the predictor uses an embedding dimensionality of 512, context window size of ten frames, and applies a residual connection from the predictor input to its output.
Text Encoder
$\text{TextOCVP}_{\text{DINO}}$ leverages a pretrained and frozen small version of T5 encoder (Raffel et al. 2020), which consists of six T5 blocks. This text encoder uses a vocabulary with size 32,128.
Scene Parsing
The scene parsing module generates $N_{\mathbf{S}}=10$ slots of dimension 128. As feature extractor, we use DINOv2 ViT-Base (Oquab et al. 2024), featuring 12 layers, using a patch size of 14, and producing patch features with dimension $D_{h}=768$. The Slot Attention corrector module processes the first video frame with three iterations in order to obtain a good initial object-centric decomposition, and a single iteration for subsequent frames, which suffices to recursively update the slot representation. The initial object slots $\textbf{S}_{0}$ are randomly sampled from a Gaussian distribution with learned mean and covariance. We use a single Transformer encoder block as the transition function, which consists of four attention heads and an MLP with a hidden dimension of 512.
Video Rendering
The video rendering module consists of two distinct decoders. First, a four-layer MLP-based Spatial Broadcast Decoder (Watters et al. 2019) with hidden dimension 1024 reconstructs the patch features from the slots. Then, a CNN-based decoder reconstructs full-resolution images from these features. It consists of four convolutional layers, each using $3\times 3$ kernels, a ReLU activation function and bilinear upsampling. A final convolutional layer and bilinear interpolation are applied to map the outputs to the RGB channels and spatial dimensions of the image.
Training
We train our model for object-centric decomposition using video sequences of length five frames for 1000 epochs. We use batch size of 16, the Adam optimizer (Kingma and Ba 2015), and a base learning rate of $4\times 10^{-4}$, which is linearly warmed-up for the first 10000 steps, followed by cosine annealing for the remaining of the training process. Moreover, we clip the gradients to a maximum norm of 0.05. The predictor module is trained given the frozen and pretrained object-centric decomposition model for 700 epochs to predict the subsequent nine frames using a single seed frame. The predictor is trained using the same hyper-parameters as for object-centric decomposition. In the predictor loss function $\mathcal{L}_{\text{TextOCVP}}$, we set $\lambda_{\text{Img}}=1$ and $\lambda_{\text{Slot}}=1$.
B.2 $\text{TextOCVP}_{\text{SAVi}}$
$\text{TextOCVP}_{\text{SAVi}}$ uses the same text-conditioned predictor and text-encoder architectures as $\text{TextOCVP}_{\text{DINO}}$, but employs SAVi (Kipf et al. 2022) as the object-centric decomposition module.
Scene Parsing
The scene parsing module generates $N_{\mathbf{S}}=8$ slots of dimension 128. Following Kipf et al. (2022), we use as feature extractor a four-layer CNN with ReLU activation function, where each convolutional layer features 32 $5\times 5$ kernels, $\text{stride}=1$, and $\text{padding}=2$. The Slot Attention corrector follows the same structure as in $\text{TextOCVP}_{\text{DINO}}$.
Video Rendering
Following Kipf et al. (2022), we utilize a CNN-based Spatial Broadcast Decoder (Watters et al. 2019) with four convolutional layers with 32 kernels of size $5\times 5$ , $\text{stride}=1$, and $\text{padding}=2$. A final convolutional layer maps from the hidden 32-channel representation to four output channels (RGB + alpha mask).
Training
We train our model for object-centric decomposition using video sequences of length ten frames for 1000 epochs, using batch size of 64, and an initial learning rate of $10^{-4}$, which is warmed up for 2500 steps, followed by cosine annealing for the remaining of the training process. Moreover, we clip the gradients to a maximum norm of 0.05. The predictor module is trained given the frozen and pretrained object-centric decomposition model for 1400 epochs to predict the subsequent nine frames using a single seed frame. The predictor is trained using the same hyper-parameters as for object-centric decomposition. In the predictor loss function $\mathcal{L}_{\text{TextOCVP}}$, we set $\lambda_{\text{Img}}=1$ and $\lambda_{\text{Slot}}=1$.
Table 4: Number of learnable parameters in TextOCVP and baselines for experiments on the CLIPort dataset.
Model # Parameters
TextOCVP 33.76M
Non-OC 34.16M
$\text{MAGE}_{\text{DINO}}$ 32.11M
SEER 405.89M
Appendix C Baselines
[…]
Table 5: Comparison of MAGE variants on CLIPort. $\text{MAGE}_{\text{DINO}}$ clearly outperforms the original MAGE variant.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
MAGE 7.116 0.453 0.713
$\text{MAGE}_{\text{DINO}}$ 23.723 0.940 0.064
[…]
MAGE is an autoregressive text-guided video prediction framework that utilizes a VQ-VAE (Van Den Oord, Vinyals et al. 2017) encoder-decoder architecture to learn efficient visual token representations. A cross-attention module aligns textual and visual embeddings to produce a spatially-aligned motion representation termed Motion Anchor (MA), which is fused with visual tokens via an axial transformer for video generation. For experiments on CATER, we use a codebook size of $512\times 256$ with a downsampling ratio of four, whereas on CLIPort we use a codebook size of $512\times 1024$.
To ensure a fair comparison with TextOCVP on CLIPort, we replace MAGE’s standard CNN encoder and decoder with the DINOv2 ViT encoder and CNN decoder used in our TextOCVP model. We refer to this modified version as $\text{MAGE}_{\text{DINO}}$. Table 5 presents a comparison between the original MAGE model and $\text{MAGE}_{\text{DINO}}$ on CLIPort. The results demonstrate that $\text{MAGE}_{\text{DINO}}$ significantly outperforms the original variant, enabling a fair comparison with TextOCVP and other baselines on this benchmark.
C.2 SEER
SEER is a diffusion-based model for language-guided video prediction. It employs an Inflated 3D U-Net derived from a pretrained text-to-image 2D latent diffusion model (Rombach et al. 2022), extending it along the temporal axis and integrating temporal attention layers to simultaneously model spatial and temporal dynamics. For the language conditioning module, SEER introduces a novel Frame Sequential Text (FSText) Decomposer, which decomposes global instructions generated by the CLIP text encoder (Radford et al. 2021) into frame-specific sub-instructions. These are aligned with frames using a transformer-based temporal network and injected into the diffusion process via cross-attention layers. We initialize SEER from a checkpoint pretrained on the Something-Something V2 dataset (Goyal et al. 2017), and further fine-tune it for a few epochs. We observed that incorporating a text loss enhanced SEER’s performance, while other hyper-parameters were kept consistent with its original implementation.
Appendix D Datasets
D.1 CATER
CATER (Girdhar and Ramanan 2020) is a dataset that consists of long video sequences, each described by a textual caption. The video scenes consist of multiple 3D geometric objects in a 2D table plane, which is split into a $6\times 6$ grid with fixed axis, allowing the exact description of object’s positions using coordinates. The text instruction describes the movement of specific objects through four atomic actions: ‘rotate’, ‘pick-place’, ‘slide’, and ‘contain’. The caption follows a template consisting of the subject, action, and an optional object or end-point coordinate, depending on the action. The movement of the objects starts at the same time step. Furthermore, the initial positions are randomly selected from the plane grid, and the camera position is fixed for every sequence.
In our work, we employ CATER-hard, which is a complete version of the CATER dataset, containing 30000 video-caption pairs, with video frames resized to $64\times 64$. It includes 5 possible objects: cone, cube, sphere, cylinder, or snitch, which is a special small object in metallic gold color, shaped like three intertwined toruses. Furthermore, every object is described by its size (small, medium, or large), material (metal or rubber), and color (red, blue, green, yellow, gray, brown, purple, cyan, or gold if the object is the snitch), and this description is included in the textual caption. Every atomic action is available. The ‘rotate’ action is afforded by cubes, cylinders and the snitch, the ‘contain’ action is only afforded by the cones, while the other two actions are afforded by every object. Every video has between 3 and 8 objects, and two actions happen to different objects at the same time. The vocabulary size is 50.
D.2 CLIPort
CLIPort (Shridhar, Manuelli, and Fox 2022) is a robot manipulation dataset, consisting of video-caption pairs, i.e. long videos whose motion is described by a textual video caption. There are many variants of the CLIPort dataset, but we focus on the Put-Block-In-Bowl variant. We generate 21000 video-caption pairs with resolution $336\times 336$. Every video contains 6 objects on a 2D table plane, and a robot arm. Objects can be either a block or a bowl, and there is at least one of them in every sequence. The starting position of each object is random, with the only constraint being that it must be placed on the table. Each video describes the action of the robot arm picking a block, and putting it in a specific bowl. The video caption follows the template ‘put the [color] block in the [color] bowl’. Each individual object in the scene has a different color. In the train and validation set, the block and the bowl that are part of the caption can have one of the following colors: blue, green, red, brown, cyan, gray, or yellow, while in the evaluation set with unseen colors they can have blue, green, red, pink, purple, white, or orange color. The other 4 objects, called distractors, can have any color. During a video sequence, it can be possible that the robot arm goes out of frame, and comes back in later frames, thus requiring the model to leverage long range dependencies. The vocabulary size is 15.
[…]
CATER1→9 CATER1→19 CLIPort1→9 CLIPort1→19
[…]
OCVP 29.08 0.874 0.078 4.16 28.11 0.854 0.101 8.08
Non-OC 29.68 0.874 0.092 3.04 28.39 0.849 0.112 8.62 23.44 0.901 0.184 8.13 20.14 0.872 0.210 13.23
SEER 22.05 0.723 0.245 11.23 16.05 0.535 0.299 17.29 21.01 0.887 0.141 6.80 11.30 0.622 0.331 8.29
MAGE 34.91 0.877 0.108 3.46 34.76 0.871 0.111 5.88 23.72 0.940 0.064 2.11 22.27 0.931 0.075 2.59
TextOCVP 32.98 0.922 0.036 2.16 31.29 0.902 0.044 5.09 26.99 0.950 0.062 1.36 23.88 0.931 0.078 2.23
Appendix E Evaluation Metrics
To measure TextOCVP’s video prediction performance and compare it with existing approaches, we evaluate the visual quality of predicted video frames using popular image- and video-based metrics.
PSNR and SSIM (Wang et al. 2004) measure pixel-wise and statistical differences between the predicted and ground-truth video frames, respectively.
LPIPS (Zhang et al. 2018) is a perceptual metric that measures the visual similarity between two images based on deep features from pretrained neural networks, usually VGG (Simonyan and Zisserman 2015). Unlike pixel-wise metrics, LPIPS compares activations from multiple layers and captures differences in texture, structure, and semantics, thus aligning closely with human perception.
While these metrics focus on frame-level visual quality, video prediction evaluation requires quantifying temporal consistency and motion realism. JEDi (Luo et al. 2025) measures the quality and realism of generated videos by comparing feature distributions of generated and real videos, capturing both visual fidelity and motion dynamics. We prefer JEDi over FVD (Unterthiner et al. 2019), as JEDi is less sensitive to small evaluation datasets, and more stable in low-motion synthetic scenarios.
In our evaluations and comparisons with baselines, we favor the LPIPS and JEDi scores, which correlate well with human perception, while reporting other metrics for completeness.
Appendix F Additional Results
F.1 SAVi vs. DINOSAUR
Current object-centric approaches for video prediction are limited to relatively simple synthetic datasets, and struggle to scale beyond scenes featuring simple 3D shapes with simple deterministic motion. We attribute this limitation primarily to the object-centric modules used for learning object representations. Motivated by this observation, we extend the recent DINOSAUR (Seitzer et al. 2023) framework to recursively process video data and reconstruct video frames from their corresponding object-centric representations.
To demonstrate the significance of the object-centric module in scaling to more complex datasets, we compare both SAVi and our Extended DINOSAUR trained on CLIPort. As illustrated in Fig. 8, SAVi struggles to accurately represent the objects on the table, missing multiple objects and changing their shape and color. In contrast, the Extended DINOSAUR model successfully reconstructs the scene, closely resembling the input, while accurately representing each object. The visual features extracted by the DINOv2 (Oquab et al. 2024) encoder contain high-level semantic information, and during training, the slots are specifically optimized to efficiently encode this information. This design enables the model to scale and handle more complex object-centric video data effectively.
$t=1$
22
33
44
55
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
GT
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
SAVi
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Ext. DINO
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Ext. DINO
Masks
Figure 8: Comparison between SAVi and our Extended DINOSAUR (Ext. DINO) decomposition modules for reconstructing a CLIPort sequence. We visualize the reconstructed frames, as well as the slot masks obtained by Extended DINOSAUR. SAVi fails to reconstruct most objects, whereas Extended DINOSAUR accurately reconstructs the scene, while representing each object.
F.2 Quantitative Comparison
In the main paper, we present quantitative evaluations on the CATER and CLIPort datasets using three evaluation metrics, shown in two separate tables.
For completeness, we report in Table 6 a text-guided video prediction evaluation of TextOCVP and multiple baseline models on CATER and CLIPort using four distinct evaluation metrics for different prediction horizons.
[…]
F.3 Computational Efficiency
In Table 7 we report the average inference time for $T=1$ and $T=9$ frame predictions on CATER and CLIPort using a NVIDIA-A6000-48Gb GPU.
[…]
Table 7: Average inference (Inf.) time for $T=1$ and $T=9$ frame predictions on CATER and CLIPort.
CATER Inf. [ms] CLIPort Inf. [ms]
Model $T=1$ $T=9$ $T=1$ $T=9$
MAGE 15.4 $\pm$ 1 111.1 $\pm$ 11 34.4 $\pm$ 3 198.8 $\pm$ 2
[…]
Table 8: Quantitative evaluation on a CLIPort test set with more objects. We report the video generation performance and its drop relative to the original dataset.
CLIPort1→9 CLIPort1→19
[…]
$\text{MAGE}_{\text{DINO}}$ 0.929 -1.2% 0.088 -37% 0.920 -1.2% 0.094 -25%
TextOCVP 0.936 -1.5% 0.076 -23% 0.921 -1.1% 0.090 -15%
F.4 Robustness to Number of Objects
We evaluate the performance of TextOCVP and $\text{MAGE}_{\text{DINO}}$ on a CLIPort evaluation set consisting of scenes with a larger number of objects than encountered during training, i.e. the 2D table plane contains 8 objects instead of the default 6 objects. This evaluation benchmarks the robustness of models when generalizing to scenes with a previously unseen number of objects. The quantitative results are presented in Table 8. We report both the video prediction performance, as well as the performance drop relative to the original evaluation set.
We observe that our model outperforms $\text{MAGE}_{\text{DINO}}$ on every metric for both prediction horizons. Notably, TextOCVP demonstrates significantly higher robustness to scenes with more objects, as reflected by the smaller decline in performance for the perceptual LPIPS metric. This result is further illustrated in Fig. 9, where we show qualitative comparisons of video generations for scenes with 8 objects. As observed, our model correctly predicts sequences following the motion described in the text instructions, whereas $\text{MAGE}_{\text{DINO}}$ fails to generate accurate sequences according to the descriptions.
These results further demonstrate the effectiveness of object-centric representations for video prediction, as TextOCVP is able to generalize to scenes with more objects by simply increasing the number of slots. This flexibility is enabled by initializing object slots via sampling from a learned Gaussian, allowing the use of a variable number of slots at test time while breaking symmetry and preserving permutation invariance. Although the object slots do not specialize during training, they reliably bind to meaningful entities through iterative attention during inference. This results in robust and scalable scene decomposition, enabling accurate modeling of complex scenes with varying object counts.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption‘put the gray block in the brown bowl Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
$\text{MAGE}_{\text{DINO}}$
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
TextOCVP
t=1t=1551010202030304040a)
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption‘put the yellow block in the gray bowl Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
$\text{MAGE}_{\text{DINO}}$
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
TextOCVP
t=1t=110102020303040405050b)
Figure 9: Qualitative evaluation of $\text{MAGE}_{\text{DINO}}$ and TextOCVP on CLIPort sequences with more objects than those seen during training. TextOCVP correctly generates sequences according to the text instructions, whereas $\text{MAGE}_{\text{DINO}}$ misses the target bowl.
Table 9: Impact of visual artifacts on TextOCVP’s performance on CLIPort. Cropping the artifact-prone bottom part of the predicted frames leads to a notable improvement in TextOCVP’s performance.
CLIPort1→9 CLIPort1→19
[…]
Full Image 0.950 0.062 0.931 0.078
Excluded Bottom 0.953 0.050 0.932 0.069
F.5 Impact of Visual Artifacts
As already discussed in the main paper, TextOCVP occasionally generates visual artifacts on the CLIPort dataset, most noticeably in the bottom-center region of the frame, where blurry patches often appear.
To assess the impact of these visual artifacts on TextOCVP’s quantitative performance, we evaluate our model after removing the last bottom rows from the predicted frames—an area that contains only background pixels and the artifacts. As shown in Table 9, removing these rows leads to improved results, particularly for the perceptual LPIPS metric. These findings verify that TextOCVP generates future frame predictions that closely follow the text description, and that its overall performance is underestimated due to localized visual artifacts. Mitigating such artifacts through improved architectures or training strategies presents an interesting direction for future work.
F.6 Text-To-Slot Attention
An additional advantage of using object-centric representations is improved interpretability. This is illustrated by visualizing the text-to-slot attention weights, which reveal how textual information guides the model’s predictions.
In Fig. 11, we visualize the text-to-slot attention weights from different cross-attention heads for a single slot that represents a rotating red cube. We observe that the slot attends to distinct relevant text tokens from the language description on each attention head, including the object shape (head 2), size (head 1) and the action taking place (head 4).
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
‘put the gray block in the brown bowl
[…]
TextOCVPt=1t=12020303040405050
Figure 10: Qualitative result on CLIPort. Top row shows ground truth frames. TextOCVP completes the pick-and-place task, whereas MAGE fails to predict the robot motion.
F.7 Additional Qualitative Evaluations
Fig. 10 shows an example where we evaluate TextOCVP and $\text{MAGE}_{\text{DINO}}$ for text-guided video prediction over a long prediction horizon of 50 frames. $\text{MAGE}_{\text{DINO}}$ fails to complete the task outlined in the textual description, as it stops generating consistent robot motion after 30 frames. In contrast, TextOCVP successfully predicts future frames where the robot completes the pick-and-place task.
Figs. 12 and 13  show qualitative evaluations on CATER in which both MAGE and TextOCVP successfully predict sequence continuations following the instructions from the textual description.
Fig. 14 illustrates an example where MAGE fails to generate a correct sequence, while TextOCVP successfully completes the task described by the text.
Figs. 15 and 16  show examples of TextOCVP’s control over the predictions. In both sequences, TextOCVP generates a correct sequence given the text instructions, and seamlessly adapts its generations to a modified version of the textual instructions.
‘the small red rubber cube is rotating.
the small gray cylinder is rotating.’
Refer to captionSceneRefer to captionAttended ObjectRefer to captionRefer to caption1.00.50.0
Head 1
the
small
red
rubber
cu
be
is
rotating
the
small
gray
metal
cylinder
is
rotating
Refer to captionRefer to caption1.00.50.0
Head 2
the
small
red
rubber
cu
be
is
rotating
the
small
gray
metal
cylinder
is
rotating
Refer to captionRefer to caption1.00.50.0
Head 3
the
small
red
rubber
cu
be
is
rotating
the
small
gray
metal
cylinder
is
rotating
Refer to captionRefer to caption
1.00.50.0
Head 4
the
small
red
rubber
cu
be
is
rotating
the
small
gray
metal
cylinder
is
rotating
Figure 11: Visualization of the text-to-slot attention weights between an object slot and the textual tokens. We visualize the text-to-slot attention weights for four different heads. Our predictor module focuses on distinct and relevant text tokens on each head, including the object size, shape or the action.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
‘the large brown metal cone is picked up and containing the
medium red rubber cone. the small gold metal snitch is rotating.’
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
MAGE
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
TextOCVP
t=1t=1551010151520203030
Figure 12: Qualitative evaluation on CATER. Both MAGE and TextOCVP successfully generate a sequence following the instructions from the textual description.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
‘the medium green rubber cone is picked up and
containing the small gold metal snitch.
the large purple rubber cone is picked up and placed to (-1, 3).’
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
MAGE
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
TextOCVP
t=1t=1551010151520203030
Figure 13: Qualitative evaluation on CATER. Both MAGE and TextOCVP successfully generate a sequence that illustrates the motion described in the text, but MAGE’s predictions are of a lower resolution.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
‘the large yellow rubber cone is sliding to (2, 3).
the small gold metal snitch is picked up and placed to (-3, 1).’
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
MAGE
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
TextOCVP
t=1t=1551010151520203030
Figure 14: Qualitative evaluation on CATER. MAGE fails to generate a sequence that accurately follows the motion described in the text. Specifically, the yellow cone does not slide as expected, and artifacts such as the merging of two small objects are introduced. On the other hand, the sequence generated by TextOCVP is closely aligned with the ground truth, accurately capturing the motion of the objects.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption‘put the blue block in the brown bowl
Original
Caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption‘put the blue block in the red bowl
Changed
Bowl
t=1t=110102020303040405050
Figure 15: Qualitative evaluation of TextOCVP controllability on CLIPort. TextOCVP correctly generates a sequence where the robot picks up and places the block specified in the textual instruction.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption‘put the green block in the cyan bowl
Original
Caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption‘put the green block in the red bowl
Changed
Bowl
t=1t=110102020303040405050
Figure 16: Qualitative evaluation of TextOCVP controllability on CLIPort. TextOCVP correctly generates a sequence where the robot picks up and places the block specified in the textual instruction.